Millard Corporation is a wholesale distributor of office products. It purchases office products from manufacturers and distributes them in the West, Central, and East regions. Each of these regions is about the same size and each has its own manager and sales staff. The company has been experiencing losses for many months. In an effort to improve performance, management has requested that the monthly income statement be segmented by sales region. The company’s first effort at preparing a segmented income statement for May is given below.

Sales Region
West Central East
Sales $311000 $805,000 $702,000
Regional expenses (traceable):
Cost of goods sold 100,000 239,000 318,000
Advertising 101,000 238,000 241,000
Salaries 54,000 57,000 106,000
Utilities 9,300 15,700 13,800
Depreciation 19,000 28,000 28,000
Shipping expense 15,000 27,000 39,000
Total regional expenses 298,300 604,700 745,800
Regional income (loss) before corporate expenses 12,700 200,300 (43,800)
Corporate expenses:      
Advertising (general) 17,000 41,000 36,000
General administrative expense 21,000 21,000 21,000
Total corporate expenses 38,000 62,000 57,000
Net operating income (loss) $(25,300) $138,300 $(100,80

The cost of goods sold and shipping expense are both variable. All other costs are fixed.

Required:

3. Prepare a new contribution format segmented income statement for May. (Round percentage answers to 1 decimal place.)
Answer #1

2 Total Company West Central East Amount % Amount % Amount Amount $ 1,818,000 100.0% $ 311,000 100.0% $ 805,000 100.0% $ 702,

A H Total Company West Amount Amount =+D3+F3+H3 =+B3/B3 311000 Central Amount =+D3/D3 805000 East Amount 702000 % =+F3/F3 =+H